OpenStudy (anonymous):

The men's world record for swimming 1500.0 m in a long course pool (as of 2007) is 14 min 34.56s. At this rate, how many seconds would it take to swim 0.350 miles? (1 mi=1609)

OpenStudy (anonymous):

14 min and 34.56 seconds is 60*14+ 34.56 = 874.56seconds. so 1500meters in 874.56 seconds. This is 1500/874.56 = 1.715m/s in velocity. Then divide the velocity with the new distance of 1609m --> 1609/1.715 = 938.11 seconds which is 15minutes and 38.11 seconds

OpenStudy (anonymous):

Would 15 minutes change to 900 seconds making the total 938.11 seconds?

OpenStudy (anonymous):

yes you devide 938.11 by 60 and then get 15.something the whole number is the minutes